OpenStudy (anonymous):

what is the domain of validity for the trigonometric identity cscθ=1/sinθ HELP PLEASE

OpenStudy (anonymous):

@jim_thompson5910 @amistre64 I've been stuck on this problem for so long and I really need the answer to this and a few others!

OpenStudy (anonymous):

I don't know @jim_thompson5910

OpenStudy (anonymous):

sin? theta? i'm really not sure

OpenStudy (anonymous):

0? I don't know. There aren't numbers in the equation..

OpenStudy (anonymous):

I don't know how to do that.. Sorry, I've been doing math homework all day and my brain is kind of fried, lol.

OpenStudy (anonymous):

I'm sorry, I don't know how to write the problem and do it in order to solve it. I suck at math.

OpenStudy (anonymous):

No

OpenStudy (anonymous):

my computer isn't loading the page

OpenStudy (anonymous):

all real numbers except multiples of pi, because that's when the denominator would equal 0 (csc = 1/sin)

OpenStudy (anonymous):

thank you :D @mangorox

OpenStudy (anonymous):

No problem @icedancerfigureskater

OpenStudy (anonymous):

Can I have help with just a few more? @mangorox

OpenStudy (anonymous):

I'll try to help

OpenStudy (anonymous):

Ok. What is the radian measure of an angle whose tangent is -0.25? My answers are: -1.33, 0.24, 2.9, and 4.95

OpenStudy (anonymous):

@mangorox

OpenStudy (anonymous):

Ok let me see...

OpenStudy (anonymous):

i'm sorry I don't know the answer to this question :(

OpenStudy (anonymous):

It's okay, do you think you can help me with a question using unit circles and inverse functions? @mangorox

OpenStudy (anonymous):

I can try, it depends on the question.

OpenStudy (anonymous):

Okay. Use the unit circle to find the inverse function value in degrees. cos^-1 (sqrt3/2)

OpenStudy (anonymous):

Does this help?: the unit circle can help you find the correct quadrants, but you have to memorize the actual degrees. You should know the sin, cos and tan of 0º, 30º, 45º, 60º, and 90º there is a pattern for sin: 0 1/2 sqrt(2)/2 sqrt(3)/2 1 if we use sqrt(0) is 0 and sqrt(1) is 1 and the sqrt(4) is 2, the pattern is sqrt(0)/2 sqrt(1)/2 sqrt(2)/2 sqrt(3)/2 sqrt(4)/2 cosine goes in the other direction tangent is sin/cos

OpenStudy (anonymous):

Kind of, but I don't know how to get my answer

OpenStudy (anonymous):

ok what is the sin(60º) ? the answer is : memorize it. See above. it is a number between -1 and 1. In this case 0.866025403... and on and on rather than type all the numbers people give the equivalent sqrt(3)/2 type sin(60 degrees)= in the google search window. It will give you that number type sqrt(3)/2= in the google search window. It will give the same number

OpenStudy (anonymous):

This might help also: http://www.khanacademy.org/math/trigonometry/v/basic-trigonometry

OpenStudy (anonymous):

the answer they want (I am assuming) is sin−1(3√2)=60º and 120º

OpenStudy (anonymous):

the answer is 60 degrees, thank you.

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

Do you know how to solve this one? @mangorox Simplify the trigonometric expression. sin^2 θ /1-cosθ

OpenStudy (anonymous):

Sorry, I don't know how to solve that one

OpenStudy (anonymous):

Ok, that's fine.

OpenStudy (anonymous):

17. Verify the pythagorean identity. 1+cot^2 θ =csc^2 θ @DLe @mangorox @jim_thompson5910 @amistre64 @dan815

OpenStudy (anonymous):

cot^2 θ = (cos^2 θ) / (sin^2 θ) so, 1 + (cos^2 θ) / (sin^2 θ) = [sin^2 θ + cos^2 θ] / (sin^2 θ) = 1 / (sin^2 θ) = csc^2 θ qed

OpenStudy (anonymous):

Thank you! @mangorox I only have one more that I need help with if you can. 18. Verify the identity. tan θ + cot θ = 1/ sinθ cosθ

OpenStudy (anonymous):

tan A + cot A =?= 1/(sinA cosA) tan A = sinA / cosA cot A = cosA / sinA So, sinA/cosA + cosA/sinA = 1/(sinAcosA) multiply both sides by sinAcosA sinAcosA (sinA/cosA) + sinAcosA(cosA/sinA) = 1 sin^A + cos^2A = 1

OpenStudy (anonymous):

Thank you so very much!!! @mangorox

OpenStudy (anonymous):

No problem! :) @icedancerfigureskater